OpenStudy (anonymous):

What is the radius of a circle if the area is 113.04 inches squared?

OpenStudy (anonymous):

sub in your area and go backwards, 113.04= pi(r^2), then solve for r

OpenStudy (gw2011):

The area of a circle is: A = (pi) (r^2) 113.04 = (pi) (r^2) r^2 = A/pi r^2 = 113.04/pi pi is approximately = 3.14159 r^2 = 113.04/3.14159 now get the square root of both sides r = 5.998 = 6 this is what the radius is approximately equal to